OpenStudy (anonymous):

Plastic parts produced by an injection-molding operation are checked for conformance to specifications. Each tool contains 12 cavities in which parts are produced, and these parts fall into a conveyor when the press opens. An inspector chooses 3 parts from among the 12 at random. Two cavities are affected by a temperature malfunction that results in parts that don't conform to specifications. a) what is the probability that the inspector finds exactly one nonconforming part? b) what is the probability that the inspector finds at least one nonconforming part? Just started class Im not good

random231 (random231):

where is your problem?

OpenStudy (anonymous):

Can anyone explain in steps? I am checking for answers but the steps do not make full sense to me because I don't understand why they do the calculations.....

OpenStudy (anonymous):

I understand that we take the factorial of 12 and then divide by 12-3 factorial and also 3 factorial....but how do I find the percentage of nonconformed parts? What other calculations what I need to do

OpenStudy (anonymous):

and why

random231 (random231):

probability of getting a nonconforming tool = 2/12=1/6 #1 prob of getting exactly 1 nonconforming part is: (1/6)*(5/6)*(5/6)*3

random231 (random231):

oh no i think im wrong! :(

OpenStudy (anonymous):

X = # of nonconforming parts S = {PPP, PPF, PFP, FPP, FFF, FFP, FPF, PFF} ( I think that's all of them) P(X=1) = 2/12 P(X>=1) = 7/8

OpenStudy (anonymous):

If my sample space is missing any, just add the missing amount to the fraction.

random231 (random231):

@lazy784 are u sure?

random231 (random231):

i think this will be bernoullis theorem!

OpenStudy (anonymous):

What I tried to do is this... After I got that the number of parts picked is (12)!/[ 3!(12-3)]! I then tried to find the nu,ber of parts that would be malfunctioned or (12)!/ (12-2)! then I divided.... would this make any sense?

random231 (random231):

@hartnn

OpenStudy (anonymous):

I got 0.1 or 1%

random231 (random231):

im getting confused on wat u r saying @elvisg1993 but let me show u wat i think!

OpenStudy (anonymous):

There won't be any factorials here. You aren't using combinations or permutations. Maybe to find out the amount of elements in the sample space, but not for the probabilities.

OpenStudy (anonymous):

http://answers.yahoo.com/question/index?qid=20100522174649AAlPUoL Does this person on yahoo answers make sense? Because he is trying through permutations I just cant follow him

OpenStudy (anonymous):

Oh i see. My sample space was woefully inadequate.

random231 (random231):

yeah he is right

OpenStudy (anonymous):

I dont understand why he gets 10C2 what does that even mean why 10

random231 (random231):

okay as there are total 10 unaffected parts he is finding the no of ways 2 unaffected parts can be selected from 10 unaffected ones!

OpenStudy (anonymous):

Why wouldn't it be 2 from 12 since there are 12 total? Thats where the confusion is

OpenStudy (anonymous):

ohhh..you are saying he is basically findin the ratio from affected to nonaffected.......and then dividing the percentage by the total number of parts.. is that what he is doing?

random231 (random231):

bullseye!

OpenStudy (anonymous):

why wouldn't the answer be found by 12C2 ? the probility that 2 of the 12 will be bad

OpenStudy (anonymous):

Are you both here? This is the only part of the answer that confuses me in his writing. ""When exactly one of these three is bad: there are 2 ways to pick the bad part and 10C2 = 45 ways to pick good """

OpenStudy (anonymous):

It's not 12 because 2 have already been chosen. 12-2=10. After 2 have been chosen, there are only 10 parts left to choose from.